OpenStudy (anonymous):

A 10-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 2 ft longer than the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace.

OpenStudy (anonymous):

10^2= x^2+(x+2)^2 == > 2x^2+4x-96=0 ===> (2x-12)(x+8)=0 ==> x=6 or x=-8( we don't cosider negative) ===> x=6 ft(vertical) ==> horizintal distance =8 ft